# Benchmarking LLMs via Uncertainty Quantification

### v2

@@ Appendix A Pseudo Code
2:Evaluation results in terms of $Acc$, $SS$, $CR$;
@@
    Employ the user-specified error rate and the calibration set to determine the conformal threshold.
@@
Table 6: The CR (%) results of LLMs with sizes ranging from 6B to 14B. The "Avg." column denotes the average performance across tasks. The small number in parentheses indicates the rank of the model on each task.
LLMs Acc (%) $\uparrow$ SS (%) $\downarrow$
@@
Appendix B Dataset Statistics
Appendix C Further Experimental Results
C.1 Detailed Results of LAC and APS
Table 1 and Table 2 have presented the average results derived from the two conformal score functions, namely, LAC and APS. In this part, we analyze the results associated with each conformal score function. The detailed results are reported in Table 6, Table 7, and Table 8.
Figure 7: Performance comparison of different versions of the Llama-2 series (7B to 70B) on various tasks.
Figure 8: Performance comparison of different versions of the Yi series (6B to 34B) on various tasks.
Figure 9: Performance comparison of different versions of the DeepSeek series (7B to 67B) on various tasks.
Figure 10: Performance comparison of different versions of the Falcon series (7B to 40B) on various tasks.
@@ C.1 Detailed Results of LAC and APS
C.2 Effects of Model Scale (Cont.)
Figures 7-10 illustrate the performance outcomes of the Llama-2 series, Yi series, DeepSeek series, and Falcon series. It is observed that while in general, increasing model size can lead to stronger performance, on some tasks, a larger model may display weaker performance. For example, on the DS task, Llama-2-70B exhibits inferior performance compared to Llama-2-13B across both accuracy and uncertainty. On the CI task, although Yi-34B demonstrates higher accuracy than Yi-6B, it shows higher uncertainty.
C.3 Effects of Instruction Finetuning (Cont.)
Figures 11-13 depict the average results across five tasks of the Yi series, DeepSeek series, and Falcon series, as well as their instruction-finetuned counterparts. Recall that for the instruction-finetuned version, two approaches are adopted to prepare the prompt input. The first approach adheres to the format of the instruction data, which is denoted as Chat-V1. The second approach employs the same prompt format as the base version and is denoted as Chat-V2. For the Yi series, it is observed that both Chat-V1 and Chat-V2 consistently yield inferior performance than the base model in terms of both Acc and SS. In contrast, for the DeepSeek series, both Chat-V1 and Chat-V2 exhibit enhanced performance in terms of Acc. Nevertheless, Chat-V1 results in greater uncertainty compared to the base model for both DeepSeek-7B and DeepSeek-70B. Chat-V2 also demonstrates higher uncertainty relative to the base model for DeepSeek-7B. For the Falcon series, it is observed that both Chat-V1 and Chat-V2 consistently lead to higher uncertainty than the base model, even though Chat-V2 achieves better performance in terms of Acc.
C.4 Effects of Error Rate
@@
C.5 Effects of Softmax Temperature
Recall that we utilize the softmax function to convert option logits generated by LLMs into probability values. These probability values are then employed by the LAC and APS score functions to estimate uncertainty. In practice, we can incorporate a temperature parameter $\tau$ into the softmax function to adjust the probability distributions (Chen et al., 2021), as shown in the following equation:
where $z=(z_{1},\dots,z_{m})\in\mathbb{R}^{m}$. A higher temperature results in a more uniform probability distribution (i.e. with higher entropy and "more random"), while a lower temperature leads to a sharper probability distribution, with one value dominating. In this part, we investigate how the temperature $\tau$ affects uncertainty estimation when using conformal prediction and perplexity, respectively. We modify the value of $\tau$ from 0.2 to 10.0 and report the results of InternLM-7B in Figure 15. Note that the temperature does not affect accuracy, so we omit results regarding accuracy.
It is observed that when using conformal prediction, we can obtain relatively stable performance in terms of both coverage rate and uncertainty (SS). However, perplexity is highly sensitive to the temperature $\tau$. When $\tau$ takes small values, perplexity results in high certainty but low coverage rate. Note that when we vary the value of $\tau$, we do not change the LLM’s accuracy. Thus, a low temperature causes the LLM to be overconfident. In this scenario, it is actually not ideal to estimate low uncertainty. Conformal prediction penalizes this phenomenon by producing large prediction sets (implying high uncertainty), which is more desirable. It is also observed that when $\tau$ takes large values and the probability distribution is close to a uniform distribution (indicating high entropy and that the LLM becomes underconfident), perplexity produces large prediction sets, which are uninformative. In contrast, conformal prediction is able to produce compact prediction sets consistently. This advantage is attributed to the use of a calibration set in conformal prediction. In summary, these observations suggest that conformal prediction is a more reliable method than entropy for evaluating uncertainty.
C.6 Unify All Tasks as One
@@
C.7 Rate of Predicted Option Being E or F
When preparing datasets, in order to enhance the complexity of tasks and effectively quantify uncertainty, two additional answer choices, E ("I don’t know") and F ("None of the above"), are incorporated into the datasets for each task. It is important to note that neither of these options represents the correct answer for any of the questions. With this in mind, our study aims to investigate whether an LLM might predict options E or F as the answer, and if so, the number of test instances for which such predictions would be made. Table 10 presents the results of various LLMs, demonstrating that, across all LLMs, only a tiny proportion of test instances are predicted to have a true answer of either E or F. Furthermore, we observe that for a particular LLM, there can be no questions whose predicted answer is E or F on some tasks (e.g., Mistral-7B on the RC task). Consequently, the inclusion of these additional options does not significantly impact the accuracy of the evaluation process. This indicates that our approach to incorporating options E and F effectively increases the difficulty of the tasks without compromising the reliability of the assessment. With more options in the answer choices, we can also quantify uncertainty more accurately.
@@ C.7 Rate of Predicted Option Being E or F
Yi-34B Base 73.19 1.40
Shared Instruction 69.87 1.46
Task-specific Instruction 71.35 1.42
Qwen-72B Base 58.30 1.70
Shared Instruction 61.08 1.67
Task-specific Instruction 62.51 1.61
Llama-2-70B Base 56.66 2.08
Shared Instruction 56.18 2.21
Task-specific Instruction 59.38 2.18
DeepSeek-67B Base 55.60 2.10
Shared Instruction 56.66 2.03
Task-specific Instruction 56.34 2.18
C.8 Comparison of Prompting Strategies
In Section 5.1, we have introduced three prompting strategies and our previous analyses are based on the average results obtained from these prompting strategies, aiming to reduce the influence of LLMs’ sensitivities to prompts. In this subsection, we delve deeper into the comparative performance of these prompting strategies. Specifically, we conduct experiments on the DS task and report the results of Yi-34B, Qwen-72B, Llama-2-70B, and DeepSeek-67B in Table 11. It can be observed that while the performance of each LLM varies with different prompting strategies, the discrepancies are generally marginal. Of greater significance is the observation that different LLMs demonstrate preferences for different prompting strategies. For example, the base prompting strategy yields the highest accuracy for Yi-34B. Conversely, Qwen-72B performs optimally with the task-specific instruction prompting strategy, while DeepSeek-67B exhibits superior accuracy with the shared instruction prompting strategy. Similar patterns can be observed from the results of the SS metric. These observations highlight the importance of considering multiple prompting strategies when benchmarking LLMs. By averaging the results from various strategies, we can mitigate the impact of prompt sensitivities and ensure a more equitable comparison of LLM performance.

### v3

@@ Appendix A Pseudo code
2:Evaluation results in terms of evaluation metrics $Acc$, $SS$, and $CR$;
@@
    Employ the user-specified error rate $\alpha$ and the calibration set to determine the conformal threshold.
@@ Appendix B Dataset statistics
Appendix C Further experimental results
C.1 Detailed results of LAC and APS
Table 6.3 has presented the average results derived from the two conformal score functions, namely, LAC and APS. In this part, we analyze the results associated with each conformal score function. The detailed results are reported in Table 7, Table 8, and Table 9.
Table 7: The CR (%) results of LLMs with sizes ranging from 6B to 14B. The "Avg." column denotes the average performance across tasks. These results correspond to using LAC and APS as the conformal score function separately.
@@ C.1 Detailed results of LAC and APS
LLMs Acc (%) $\uparrow$ SS $\downarrow$
@@
Figure 6: Performance comparison of different versions of the Llama-2 series (7B to 70B).
Figure 7: Performance comparison of different versions of the Yi series (6B to 34B).
Figure 8: Performance comparison of different versions of the DeepSeek series (7B to 67B).
Figure 9: Performance comparison of different versions of the Falcon series (7B to 40B).
@@
C.2 Effects of model scale (cont.)
Figures 6-9 illustrate the performance outcomes of the Llama-2 series, Yi series, DeepSeek series, and Falcon series. It is observed that while in general, increasing model size can lead to stronger performance, on some tasks, a larger model may display weaker performance. For example, on the DS task, Llama-2-70B exhibits inferior performance compared to Llama-2-13B across both accuracy and uncertainty. On the CI task, although Yi-34B demonstrates higher accuracy than Yi-6B, it shows higher uncertainty.
C.3 Effects of instruction finetuning (cont.)
Figures 10-12 depict the average results across five tasks of the Yi series, DeepSeek series, and Falcon series, as well as their instruction-finetuned counterparts. Recall that for the instruction-finetuned version, two approaches are adopted to prepare the prompt input. The first approach adheres to the format of the instruction data, which is denoted as Chat-V1. The second approach employs the same prompt format as the base version and is denoted as Chat-V2. For the Yi series, it is observed that both Chat-V1 and Chat-V2 consistently yield inferior performance than the base model in terms of both Acc and SS. In contrast, for the DeepSeek series, both Chat-V1 and Chat-V2 exhibit enhanced performance in terms of Acc. Nevertheless, Chat-V1 results in greater uncertainty compared to the base model for both DeepSeek-7B and DeepSeek-70B. Chat-V2 also demonstrates higher uncertainty relative to the base model for DeepSeek-7B. For the Falcon series, it is observed that both Chat-V1 and Chat-V2 consistently lead to higher uncertainty than the base model, even though Chat-V2 achieves better performance in terms of Acc.
LLMs CR (%) Acc (%) $\uparrow$ SS $\downarrow$
Mixtral-8x7B 92.59 73.74 2.03
Mistral-7B 91.78 64.14 2.43
Llama-2-13B 91.60 60.42 2.56
Qwen-14B 92.94 74.00 2.17
C.4 Effects of mixture of experts
A Mixture of Experts (MoE) is a technique that combines multiple specialized models with a gating mechanism to enhance performance by leveraging diverse expertise and adaptability in handling complex data relationships. In recent months, the adoption of the MoE technique to augment the performance of LLMs has been steadily increasing. Considering this, we study how MoE impacts the uncertainty of LLMs, specifically comparing the Mixtral-8x7B MoE model [^6] https://huggingface.co/mistralai/Mixtral-8x7B-v0.1 with Mistral-7B. The results are reported in Table 10. While Mixtral-8x7B has 46.7B total parameters, it only uses 12.9B parameters per token. It, therefore, processes input and generates output at the same speed and for the same cost as a 12.9B model. [^7] https://mistral.ai/news/mixtral-of-experts/ To provide a comprehensive comparison, we also include the results of Llama-2-13B and Qwen-14B. As can be observed, Mixtral-8x7B consistently outperforms Mistral-7B across both accuracy and uncertainty. Remarkably, it achieves accuracy levels comparable to Qwen-14B, while demonstrating the lowest average uncertainty among the four LLMs. These observations show that MoE is indeed an effective method to enhance the accuracy and reduce the uncertainty of LLMs.
C.5 Effects of error rate
@@
C.6 Effects of amount of calibration data
As described in § 3, conformal prediction requires a calibration set to calculate the threshold $\hat{q}$. In our prior analyses, we have allocated 50% of the data as the calibration set and the remaining 50% as the test set. Here, we explore the impact of varying the proportion of calibration data, ranging from 10% to 50%, on uncertainty quantification. Note that the same 50% of data is consistently used as the test set. The average results across five tasks of the Llama-2 series are shown in Figure 14. It is observed that there are no significant variations in coverage rate (CR) and uncertainty (SS) for all versions of the Llama-2 series when varying the amount of calibration data. This observation confirms the efficacy of applying conformal prediction for uncertainty quantification in our analysis.
C.7 Effects of softmax temperature
Recall that we utilize the softmax function to convert option logits generated by LLMs into probability values. These probability values are then employed by the LAC and APS score functions to estimate uncertainty. In practice, we can incorporate a temperature parameter $\tau$ into the softmax function to adjust the probability distributions [11], as shown in the following equation:
where $z=(z_{1},\dots,z_{m})\in\mathbb{R}^{m}$. A higher temperature results in a more uniform probability distribution (i.e. with higher entropy and "more random"), while a lower temperature leads to a sharper probability distribution, with one value dominating. In this part, we investigate how the temperature $\tau$ affects uncertainty estimation when using conformal prediction and perplexity, respectively. We modify the value of $\tau$ from 0.2 to 10.0 and report the results of InternLM-7B in Figure 15. Note that the temperature does not affect accuracy, so we omit results regarding accuracy and only provide results of CR and SS.
@@ C.7 Effects of softmax temperature
It is observed that when using conformal prediction, we can obtain relatively stable performance in terms of both coverage rate and uncertainty (measured by SS). However, perplexity is highly sensitive to the temperature $\tau$. When $\tau$ takes small values, perplexity results in high certainty but low coverage rate. Note that when we vary the value of $\tau$, we do not change the LLM’s accuracy. Thus, a low temperature causes the LLM to be overconfident. In this scenario, it is actually not ideal to estimate low uncertainty. Conformal prediction penalizes this phenomenon by producing large prediction sets (implying high uncertainty), which is more desirable. It is also observed that when $\tau$ takes large values and the probability distribution is close to a uniform distribution (indicating high entropy and that the LLM becomes underconfident), perplexity produces large prediction sets, which are uninformative. In contrast, conformal prediction is able to produce compact prediction sets consistently. This advantage is attributed to the use of a calibration set in conformal prediction. In summary, these observations suggest that conformal prediction is a more reliable method than entropy or perplexity for evaluating uncertainty.
@@
C.8 Unify all tasks as one
@@
C.9 Rate of predicted options being E or F
When preparing datasets, in order to enhance the complexity of tasks and effectively quantify uncertainty, two additional answer choices, E ("I don’t know") and F ("None of the above"), are incorporated into the datasets for each task. It is important to note that neither of these options represents the correct answer for any of the questions. With this in mind, our study aims to investigate whether an LLM might predict options E or F as the answer, and if so, the number of test instances for which such predictions would be made. Table 12 presents the results of various LLMs, demonstrating that, across all LLMs, only a tiny proportion of test instances are predicted to have a true answer of either E or F. Furthermore, we observe that for a particular LLM, there can be no questions whose predicted answer is E or F on some tasks (e.g., Mistral-7B on the RC task).
We also report the average prediction set size (SS) when options E and F are excluded from the answer choices to evaluate their impact on uncertainty quantification. The results, presented in Table 13, indicate that the SS values remain relatively consistent with those observed when options E and F are included. Furthermore, while smaller average SS values are usually obtained when LLMs demonstrate strong performance, larger SS values are observed even with fewer answer choices (i.e. without options E and F) when LLMs exhibit weaker performance (e.g., 3.57 vs. 3.74 for MPT-7B and 3.48 vs. 3.55 for Falcon-40B). It is also noted that Llama-2-70B, DeepSeek-67B and Yi-6B achieve the same average SS value of 1.89 when options E and F are excluded, making them not differentiable in terms of prediction uncertainty.
Consequently, the inclusion of these additional options does not significantly impact the accuracy of the evaluation process. This indicates that our approach to incorporating options E and F effectively increases the difficulty of the tasks without compromising the reliability of the assessment. With more options in the answer choices, we can also quantify uncertainty more accurately.
Table 13: The average prediction set size (SS) when the option E ("I don’t know") and option F ("None of the above") are included in or excluded from the answer choices. Note that neither of them corresponds to the ground truth answer.
LLMs With Options E and F Without Options E and F
QA RC CI DRS DS Avg. QA RC CI DRS DS Avg.
Yi-34B 2.60 1.71 1.90 1.77 1.69 1.93 2.06 1.42 1.52 1.49 1.58 1.61
Qwen-72B 2.45 1.90 1.80 2.09 2.06 2.06 2.02 1.52 1.53 1.59 1.78 1.69
Qwen-14B 2.80 1.74 2.02 1.94 2.37 2.17 2.39 1.41 1.54 1.67 2.04 1.81
Llama-2-70B 2.62 1.78 1.82 2.34 2.25 2.16 2.30 1.51 1.69 1.88 2.07 1.89
DeepSeek-67B 2.65 1.54 2.43 1.89 2.25 2.15 2.21 1.40 2.11 1.68 2.05 1.89
Yi-6B 3.20 1.92 1.88 2.85 1.96 2.36 2.41 1.54 1.77 1.99 1.76 1.89
Gemma-7B 2.72 1.88 2.04 2.14 3.11 2.38 2.42 1.69 1.97 2.05 2.96 2.22
Mistral-7B 2.80 1.75 2.48 2.71 2.40 2.43 2.48 1.67 2.48 2.53 2.28 2.29
Llama-2-13B 3.06 2.24 2.72 2.55 2.24 2.56 2.92 2.00 2.69 2.50 2.18 2.46
Qwen-7B 3.26 2.15 2.28 2.51 2.92 2.63 2.78 1.72 2.15 2.16 2.84 2.33
InternLM-7B 3.49 2.19 3.28 3.63 4.47 3.41 3.34 2.08 3.47 3.13 3.69 3.14
Llama-2-7B 3.20 2.39 3.27 3.26 3.30 3.09 3.45 2.36 3.57 3.57 2.93 3.18
DeepSeek-7B 3.34 2.77 3.06 3.40 3.08 3.13 3.41 2.42 3.45 3.61 3.23 3.22
Qwen-1.8B 3.20 2.58 3.49 3.45 4.18 3.38 3.43 2.36 3.59 3.46 3.71 3.31
Falcon-40B 3.25 3.12 3.54 3.59 3.89 3.48 3.46 3.20 3.70 3.72 3.66 3.55
MPT-7B 3.53 3.46 3.60 3.59 3.66 3.57 3.73 3.68 3.75 3.75 3.78 3.74
Falcon-7B 3.90 3.60 3.66 3.64 3.92 3.75 3.76 3.76 3.76 3.76 3.77 3.76
@@ C.9 Rate of predicted options being E or F
      Yi-34B       Base Prompt       73.19       1.40
      Shared Instruction Prompt       69.87       1.46
      Task-specific Instruction Prompt       71.35       1.42
      Qwen-72B       Base Prompt       58.30       1.70
      Shared Instruction Prompt       61.08       1.67
      Task-specific Instruction Prompt       62.51       1.61
      Llama-2-70B       Base Prompt       56.66       2.08
      Shared Instruction Prompt       56.18       2.21
      Task-specific Instruction Prompt       59.38       2.18
      DeepSeek-67B       Base Prompt       55.60       2.10
      Shared Instruction Prompt       56.66       2.03
      Task-specific Instruction Prompt       56.34       2.18
C.10 Comparison of prompting strategies
In § 5, we have introduced three prompting strategies and our previous analyses are based on the average results obtained from these prompting strategies, aiming to reduce the influence of LLMs’ sensitivities to prompts. In this subsection, we delve deeper into the comparative performance of these prompting strategies. Specifically, we conduct experiments on the DS task and report the results of Yi-34B, Qwen-72B, Llama-2-70B, and DeepSeek-67B in Table 14. It can be observed that while the performance of each LLM varies with different prompting strategies, the discrepancies are generally marginal. Of greater significance is the observation that different LLMs demonstrate preferences for different prompting strategies. For example, the base prompting strategy yields the highest accuracy for Yi-34B. Conversely, Qwen-72B performs optimally with the task-specific instruction prompting strategy, while DeepSeek-67B exhibits superior accuracy with the shared instruction prompting strategy. Similar patterns can be observed from the results of the SS metric. These observations highlight the importance of considering multiple prompting strategies when benchmarking LLMs. By averaging the results from various strategies, we can mitigate the impact of prompt sensitivities and ensure a more equitable comparison of LLM performance.
LLMs CR (%) Acc (%) $\uparrow$ SS $\downarrow$
Yi-34B 94.22 81.46 1.93
Yi-34B 93.31 80.10 1.88
Llama-2-70B 93.11 72.24 2.16
Llama-2-70B 91.55 63.55 2.72
C.11 Ablation study
Inspired by in-context learning [66], our previous analyses have incorporated demonstrations for each task. Here, we aim to compare the performance of LLMs when demonstrations are included or excluded from the prompt input. The results of Yi-34B and Llama-2-70B are reported in Table 15. We observe that the presence of demonstrations leads to improved performance for both models, as evidenced by increased accuracy. Having demonstrations also leads to lower uncertainty for Llama-2-70B but higher uncertainty for Yi-34B. Overall, these results substantiate the effectiveness of incorporating demonstrations into the prompt to stimulate the capabilities of LLMs.
C.12 Case study
@@ Appendix D Prompt templates
Appendix E Limitations
Despite the multiple advantages of conformal prediction over other uncertainty quantification methods, it demonstrates three key limitations when employed to assess the uncertainty of LLMs. Firstly, the application of conformal prediction necessitates access to model output logits, which precludes the possibility of benchmarking LLMs such as ChatGPT that are only accessible via their APIs. Secondly, the adoption of conformal prediction poses challenges in evaluating the generative capabilities of LLMs. In our analyses, all tasks are transformed into multiple-choice questions, thereby primarily assessing the language understanding abilities of LLMs rather than their generative potential. Thirdly, the prediction sets generated by conformal prediction could be significantly influenced by the conformal score function utilized. Thus, for a specific LLM, varying conformal score functions may yield disparate estimations of uncertainty. It is worth mentioning that while we have extended our benchmarking to closed-source LLMs and free-form text generation, the extension can only provide an approximation.
Nevertheless, the limitations of other uncertainty quantification methods prevent us from applying them in the context of LLMs. Per our knowledge, conformal prediction is currently the most feasible technique for robust uncertainty quantification of LLMs. Furthermore, some recent studies have tried to incorporate conformal prediction into the language generation process [55, 53, 16]. We posit that conformal prediction will eventually evolve into an appropriate method for quantifying the uncertainty of language generation in the future.
Last but not least, it is important to note that the scope of this study is limited to evaluating the capabilities of LLMs in the context of language processing exclusively. The current trend in the field is towards the development of multi-modal foundation models [44, 41, 47], which have the capacity to process multiple modalities rather than just language. Therefore, it would be a significant extension of this research to investigate the uncertainty associated with these foundation models when they are applied to non-language modalities, which constitutes an important avenue for future research.
Appendix F Societal Impacts
While benchmarking LLMs via uncertainty quantification can lead to more accurate assessments of their performance, it is crucial to consider and address any potential negative societal impacts. We list several possible concerns below:
  • Misuse of Technology: Enhanced performance and reliability of LLMs could lead to their misuse in generating misleading or harmful content, such as deepfakes, disinformation, or automated trolling. Improved uncertainty quantification might make these models more convincing and harder to detect.
  • Bias and Fairness: Even with improved uncertainty quantification, LLMs can still perpetuate and amplify existing biases present in the training data. This can lead to unfair treatment of certain groups or individuals, reinforcing stereotypes and discrimination.
  • Job Displacement: As LLMs become more capable, there is a potential for job displacement in fields that rely heavily on language processing, such as customer service, translation, and content creation. This could lead to economic and social challenges for affected workers.